OpenStudy (anonymous):

Differential Equations - just wanted to verify my work. Solve the following IVP ; (3x^2+y-2)dx + (x+2y)dy = 0 y(2) = 3 My = 1 Nx = 1 f(x,y) = Integral(M(dx) + g(y)) = Integral(3x^2 + y - 2)dx + g(y) = x^3 - 2x + g'(y) x^3 - 2x + g'(y) = x + 2y x^3 + g'(y) = 3x + 2y g'(y) = Integral(3x + 2y - x^3)dy g(y) = y^2 f(x,y) = x^3 -2x + g(y) C = x^3 - 2x + y^2 y(2) = 3 C = (2)^3 - 2(2) + (3)^2 C = 8 - 4 + 9 C = 13 13 = x^3 - 2x + y^2 0 = x^3 -2x +y^2 -13 Just wanted to verify that all of the steps I took were correct! Thanks!

OpenStudy (anonymous):

Integral(3x^2 + y - 2)dx + g(y) = x^3 + xy -2x + g(y)

OpenStudy (anonymous):

and totaly don't understand this: x^3 - 2x + g'(y) = x + 2y if you take derivative respect to y that's not what you get, :)

OpenStudy (anonymous):

you there? @Archetype

OpenStudy (anonymous):

Yes. One second.

OpenStudy (anonymous):

If you whant i explain you in detail this type of ODE

OpenStudy (anonymous):

I believe I left off the integral sign? Everything you typed looked fine to me... I didn't take the derivative with respect to y.

OpenStudy (anonymous):

= Integral(3x^2 + y - 2)dx + g(y) Derivative with respect to X, which gives this; = x^3 - 2x + g'(y)

OpenStudy (anonymous):

then you wrong, my friend ..:)

OpenStudy (anonymous):

How so? Where did I go wrong? Guess I do need an explanation.

OpenStudy (anonymous):

If I take the integral of X in the equation (x + y) then Y is treated as a constant and is = 0.

OpenStudy (anonymous):

Ok. This type of ODE is colled exact differ. eq.. It means that the equation is actualy a full differential of some function F(x,y), whose partial derivatives respect to x, Mx, and y, My are whats stands by dx and dy in the equation respectivly. Till now, ok?

OpenStudy (anonymous):

To actualy check that we got this type of equation is the step of checking the second derivatives are equal... That was your 1º step

OpenStudy (anonymous):

My=Nx

OpenStudy (anonymous):

Yeah.. I know the basics of it. I believe my work should show that. If I am making a mistake it should be something fairly simple.

OpenStudy (anonymous):

not really, :)

OpenStudy (anonymous):

Integral(3x^2 + y - 2)dx + g(y) = x^3 + xy -2x + g(y) From what i have read; that is not correct.

OpenStudy (anonymous):

Unless I'm just out to lunch. I imagine you have far more experience with these then I do.

OpenStudy (anonymous):

just keep in mind that M is partial derivative of F respect to x. And N is Fy

OpenStudy (anonymous):

So to find F you can integrate Fx, which you call M respect to x. So lets do it: Fx = 3x^2 + y - 2 keep in mind that now we treat y like a constant parameter. \[F = \int\limits (3x ^{2} +y -2) dx= x ^{3}+xy -2x +g(y)\] now, if we would take derivative respect to y of this result it should be equal to N. So let's do it: x+g'(y) = N = x+2y so g'(y) = 2y g=y^2 +C To find C use the information y(2) = 3 Ok?

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

Yes.. but it looks like my only mistake was integrating 3x^2 + y -2 incorrectly?

OpenStudy (anonymous):

notonly, you didn't take derivative respect to y, you did respect to x....

OpenStudy (anonymous):

So this, for example... Integral of (7x-3y)dx = (7/2)x^2 + 3yx ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ugh, ok, excellent. Well thank you very much for the help. I've made that mistake on every exact DE problem I've done so far.

OpenStudy (anonymous):

glad that could help

OpenStudy (anonymous):

Ugh, yeah, just typed it into Wolframalpha and it gave the same (not that i doubted you..) Thanks again mate, you were a GREAT help!

OpenStudy (anonymous):

yw